OpenStudy (anonymous):

If b is the harmonic mean between a and c, prove that \[\frac{1}{b-a}+\frac{1}{b-c}=\frac{1}{a}+\frac{1}{c}\]

mathslover (mathslover):

hmn wait lemme open my book .. wait

OpenStudy (anonymous):

If am not wrong this must be the meaning:- Harmonic Mean of a set of numbers is the number of items divided by the sum of the reciprocals of the numbers. Hence, the Harmonic Mean of a set of n numbers

OpenStudy (anonymous):

i am confused by your definition. i know that H.M. is the reciprocal of A.M.

OpenStudy (anonymous):

SO you can proceed further @shubham.bagrecha as Harmonic Mean of a set of numbers is the number of items divided by the sum of the reciprocals of the numbers.

OpenStudy (anonymous):

So hence proven

OpenStudy (anonymous):

how ??

mathslover (mathslover):

basic formula : \[\large{\textbf{Harmonic Mean}=\frac{2ab}{a+b}}\]

hartnn (hartnn):

lets start with LHS \[\frac{1}{b-a}+\frac{1}{b-c}=\frac{2b-(a+c)}{b ^{2}+ac-b(a+b)}\] now since they are in HP, \[a+c=\frac{2ac}{b}\] substituting this,we get \[\frac{2b-\frac{2ac}{b}}{b ^{2}+ac-2ac}=\frac{2}{b}\] and accordind to defination of HP,u get that as 1/c+1/a.....

mathslover (mathslover):

the better formula for this case will be : let the numbers be a, b and c so we have : \[\large{\frac{1}{b}-\frac{1}{a}=\frac{1}{c}-\frac{1}{b}}\] \[\large{\frac{2}{b}=\frac{a+c}{ac}}\] \[\large{b=\frac{2ac}{a+c}}\]

hartnn (hartnn):

^^

mathslover (mathslover):

good work @hartnn